# Human-AI Collaboration with Misaligned Preferences

### v1

@@ 6 Empirical Study
We use the sushi preference dataset (Kamishima, 2003), which consists of about 5k rankings of ten kinds of sushi, where we denote the ten types of sushi by $x_{1},\dots,x_{10}$. For computational efficiency, we consider only the partial rankings of the humans in our sushi dataset over the items $x_{1},\ldots,x_{5}$, treating these partial rankings as their ground-truth preferences. Each ranking is also associated with the fraction of the population in the dataset who had this preference over the types of sushi. A detailed breakdown of the most frequent sushi rankings is presented in Table˜3 (Appendix˜D).
Note that in this dataset, it is inherently impossible to distinguish between a) population-level heterogeneity in preferences, and b) user-level noise in reporting those preferences. To our knowledge, there is not a preference dataset that cleanly displays both a) and b). As such, in this section we make the assumption that users have reported their sushi preferences perfectly (no errors from b) and that the preference dataset simply reflects user heterogeneity (only difference is from a). In order to model user-level noise, we add noise (e.g., Mallows) on top of the ground-truth preferences within (Kamishima, 2003). Furthermore, because this dataset does not include explicit human valuation scores for the sushi items, we assume item values within each ranking is determined by their Borda counts. Formally, given a ranking $\pi$ over $m$ items $\{x_{1},\ldots,x_{m}\}$, where $\pi(i)$ denotes the position of item $x_{i}$ (with $\pi(i)=1$ being the most preferred), the Borda score assigned to item $x_{i}$ is
@@ 8.1 Model Extensions
Other extensions could change our model of human-algorithm interaction. For example, our model of subset curation assumes that the human observes the set of presented items as an unordered set and is not biased by any inherent ordering of items. This is almost certainly not true in practice, as it has been demonstrated that humans are often biased by the order in which items are presented (e.g., see (Joachims et al., 2007; Agichtein et al., 2006; Craswell et al., 2008)), and this effect has been studied in human-algorithm collaboration more specifically (e.g. (McLaughlin and Spiess, 2024; Donahue et al., 2024; Wang and Pananjady, 2022)). It could also be possible that the human and algorithm undergo a sequence of back-and-forth interactions which ultimately end up at a final solution (as in (Collina et al., 2025a)), which may lead to different implications on personalization. Finally, our results on optimizing the algorithm for different social welfare functions assume that the algorithm has perfect knowledge of the distribution of human preferences. In practice, this is almost certainly not the case, and it could be useful to study settings where the algorithm must learn human preferences over multiple interactions.
@@ 8.2 Implications for Policy
In some real-life settings, humans have near-perfect ability to identify the “correct” outcome (e.g., the political opinion that they most closely identify with). However, in many real-life settings, humans may only be able to imperfectly (noisily) recover their own true preferences: prior research has shown that humans noisily pick the favorite item from sets, such as presented movies or items on a menu (Agranov and Ortoleva, 2017; Plackett, 1975; Luce et al., 1959; Hey and Orme, 1994). In this setting, our results show that having algorithms that are “misaligned” in specific ways can be beneficial. This implies that designers of algorithms should maybe not always focus on designing algorithms that personalize to individual humans (eliminating misalignment): in settings where humans are noisy, personalizing may end up being harmful. This implies that algorithmic designers may want to take different strategies towards alignment depending on the context in which humans will use the tool, focusing on alignment more in noiseless settings and less in settings where humans are expected to know their own preferences less well.
@@
Our results also show that there can be benefits to algorithmic noise (e.g., positive temperature) when a single algorithm is serving a population of heterogeneous users. Benefits of randomization for goals related to fairness have been shown in other settings (e.g., (Jain et al., 2024; Singh and Joachims, 2018)), though to our knowledge, this question is less well studied in the context of complementarity or similar benchmarks. Most generative AI tools have a constant temperature across multiple queries: our work suggests that for certain welfare objectives, it could be helpful to have temperature change dynamically, being lower in settings where human opinions are roughly in agreement and larger in settings where human opinions are more diverse.
@@ Acknowledgements
The research of Bhaskar Ray Chaudhury and Jiaxin Song was supported by NSF CAREER Grant CCF-2441580. The research of Kate Donahue was supported by the MIT METEOR Fellowship. The research of Parnian Shahkar was supported by NSF Grant CCF-2454115. We thank Professor Ariel Procaccia (Harvard University), Professor Jann Spiess (Stanford), Kostas Kollias (Google Research), Professor Manish Raghavan (MIT), Professor Biaoshuai Tao (SJTU), as well as attendees at the Themis Science Talk at Amazon, the Human-AI workshop at EC’25, the Human-AI Complementarity for Decision Making Workshop at CMU, and the AI and Science Group at Meta, for their valuable discussions and insightful suggestions. We also thank the anonymous reviewers for their constructive feedback.
@@ A.1 Basic Properties
Let $\mathfrak{S}(M)$ be the set of all rankings of $M$. Denote by $x_{i}\succ_{\pi}S$ for a set of items $S$ if $x_{i}$ is before any item $x_{j}\in S\setminus\{x_{i}\}$. A swap $(x_{i}x_{j})$ is valid with respect to $\pi$ and a ground-truth ranking $\pi^{*}$ if $\pi$ and $\pi^{*}$ differ on the relative ranking of $x_{i}$ and $x_{j}$. By the definition of inversion-monotonicity, if $(x_{i}x_{j})$ is a valid swap with respect to any given ranking $\pi$ and a ground-truth ranking $\pi^{*}$, then $\pi\circ(x_{i},x_{j})$ has a higher probability than $\pi$. Moreover, according to [Donahue et al., 2024, Lemma 1], the number of inversions reduces by at least one after the swap.
@@ A.2 Properties of Mallows Model
$x_{i}$ ranked the first in $\pi$ $\frac{\exp(-\phi\cdot(i-1))}{Z_{m}(\phi)}$ [Awasthi et al., 2014]
$x_{1}$ ranked at the $i$-th location in $\pi$ $\frac{\exp(-\phi\cdot(i-1))}{Z_{m}(\phi)}$ [Awasthi et al., 2014]
$x_{1}$ ranked at the first $k$ locations in $\pi$ $\frac{Z_{k}(\phi)}{Z_{m}(\phi)}$ [Awasthi et al., 2014]
$x_{i}$ ranked before $x_{j}$ with $i<j$ and k=ji+1k=j-i+1 in $\pi$ $\frac{k}{1-\exp(-\phi\cdot k)}-\frac{k-1}{1-\exp(-\phi\cdot(k-1))}$ [Boehmer et al., 2023]
@@ A.2.1 Item Location
Then applying Equation˜3, since the sum of $\exp(-\phi\cdot\mathrm{inv}(\pi_{S}))$ over all permutations of $S$ is equal to $\prod_{i=1}^{k}Z_{i}(\phi)$ [Mallows, 1957], we then have
@@ A.2.2 Itemwise Comparison
Lemma 12 (Proposition 3.8 of [Boehmer et al., 2023]).
@@
We first prove it holds for the case of $i=i^{\prime}$. Hence, by assumption, $j^{\prime}\geq j$. For any ranking $\pi$ that ranks $x_{i}$ before $x_{j}$ while ranking $x_{i}$ after $x_{j^{\prime}}$, we map it to a new one by swapping $x_{j}$ and $x_{j^{\prime}}$. By [Donahue et al., 2024, Lemma 1], the number of inversions reduces by at least one, and the probability of the new ranking is weakly larger than the original one. Therefore, it implies that the probability of $x_{i}$ being before $x_{j}$ is smaller. By Lemma˜12, the probability of pairwise comparison only depends on the difference of the two indices. Therefore, the lemma holds for every $i,j,i^{\prime}$, and $j^{\prime}$. ∎
@@
We partition the set of permutations into $k$ sets based on the relative position of $x_{i(1)}$. Let $\mathfrak{S}_{t}$ be the set of permutations where $x_{i(1)}$ is placed at the $t$-th relative position among all items of $\mathcal{S}$. For any $\pi\in\mathfrak{S}_{t+1}$ with $t\leq k-1$, we map it to a permutation $\pi\in\mathfrak{S}_{t}$ by swapping $x_{i(1)}$ and the item that is placed at the $t$-th position among all items of $\mathcal{S}$. It is a valid swapping with respect to $\pi$ and $\pi^{*}$ by definition. By [Donahue et al., 2024, Lemma 1], the probability of $\pi^{\prime}$ is at least $\exp(\phi)$ times the probability of $\pi$. By summing the inequality up over all permutations of $\mathfrak{S}_{t+1}$, we can conclude that $\operatorname*{\mathbb{P}}[\pi\in\mathfrak{S}_{t}]\geq\exp(\phi)\cdot\operatorname*{\mathbb{P}}[\pi\in\mathfrak{S}_{t+1}]$. Hence, by multiplying these inequalities, we can have
@@ C.2 Mixed Integer Linear Programming Formulation
Finally, we adapt the MIP of Désir et al. [2016] and formulate the problem of finding the welfare-maximizing strategy as an MIP. Désir et al. [2016] studies the assortment optimization problem under the Mallows model, where each item is associated with a nonnegative value $r_{i}$ and the goal is to pick a subset of items $\mathcal{S}$ to maximize the weighted sum of choice probabilities. Formally, the problem can be expressed as follows:
@@
Désir et al. [2016] shows the choice probability of $x_{i}$ being the first among all items in $\mathcal{S}$ can be calculated by a dynamic programming algorithm. However, it is worth noting that the algorithm relies on prior knowledge that some item is known to be included in $\mathcal{S}$, which is further used as a “guess” of the MIP formulation. We next provide a dynamic programming of the choice probability with marginal change to the original algorithm of Désir et al. [2016] without any prior knowledge.
Let $W(i,s,t)$ be the probability of item $x_{i}$ being chosen at position $s$ after $t$ steps of repeatedly insertion, where $i\in[m]$, $s\in[t]$, and $t\in[m]$. The following algorithm assumes the ground-truth ranking $\pi^{*}=(x_{1},\ldots,x_{m})$ without loss of generality. We use the same notations for constants $p_{t+1,s}$, $\gamma_{t+1,s}$ as in Désir et al. [2016]. Probability $p_{t+1,s}$ is the probability of inserting item $x_{t+1}$ at position $s$ into a partial permutation of $x_{1},\ldots,x_{t}$. Also, $\gamma_{t,s}=\sum_{\ell=1}^{s}p_{t,\ell}$. Both sets of constants have closed-form expressions once the ground-truth ranking and the accuracy parameter are given.
@@
Compared to the original dynamic programming of Désir et al. [2016], we do not require a precondition that an item $x_{i}$ needs to be included in $\mathcal{S}$. Hence, we update the state space starting from step $t=1$ rather than $t=2$. As a result, if $x_{t}\in\mathcal{S}$, at the $t$-th step, there are two possible cases for item $x_{t}$ being selected at position $s$: 1) none of items $x_{1},\ldots,x_{t-1}$ is included in $\mathcal{S}$; therefore, item $x_{t}$ is selected at position $s$ with probability $p_{t,s}$; 2) at least one of items $x_{1},\ldots,x_{t-1}$ is included in $\mathcal{S}$; therefore, item $x_{t}$ is chosen only if all items of $x_{1},\ldots,x_{t-1}$ that are included in $\mathcal{S}$ are ranked after $x_{t}$ and $x_{t}$ is inserted at position $s$ (the same argument of Désir et al. [2016]). Otherwise, if $x_{t}\notin\mathcal{S}$, the same argument of Désir et al. [2016] still applies for updating the probabilities of choosing other items. ∎
@@ C.3 Uplift under Special Cases
Fix a set $S$ of size $k$. Notice that, when $x_{j}\notin S$, $\operatorname*{\mathbb{P}}[x_{H}=x_{j}\mid\rho[:k]=\mathcal{S}]=0$. Next, we consider the sum of the conditional probabilities for $x_{j}\in S$. We first show that, conditioned on $\rho[:k]=\mathcal{S}$, the partial ordering of the first $k$ items (denoted by $\operatorname{\rho}_{|\mathcal{S}}$) forms a Mallows distribution of items of $\mathcal{S}$. Observe that the number of inversions between $S$ and $M\setminus S$ for every permutation $\operatorname{\rho}$ such that $\rho[:k]=\mathcal{S}$ since the two sets of items are isolated. Hence, among all the permutations such that $\rho[:k]=\mathcal{S}$, the probability of each of them is proportional to $\exp(-\phi)$ raised to the power of the number of inversions inside $\mathcal{S}$, which implies that the partial ranking $\pi_{|S}$ forms a Mallows distribution of items of $\mathcal{S}$. Denote the $s$ items by $\mathcal{S}=\{x_{i(1)},\ldots,x_{i(k)}\}$ with $i(1)<\ldots<i(k)$. By the known property of Mallows model [Awasthi et al., 2014, Lemma 2.5], the probability of item $x_{i(j)}$ being the first is equal to $\exp(-\phi\cdot(j-1))/Z_{k}(\phi)$. Summing it up for $j$ from $1$ to $i$, we get the probability of $x_{H}$ being one of $x_{1}$ to $x_{i}$ is equal to $Z_{s}(\phi)/Z_{k}(\phi)$, where $s$ is the size of the intersection of $\{x_{1},\ldots,x_{i}\}$ and $\mathcal{S}$. Therefore, the probability of the human picking one of the first $i$ items is equal to

### v2

@@ 6 Empirical Study
We use the sushi preference dataset (kamishima2003nantonac), which consists of about 5k rankings of ten kinds of sushi, where we denote the ten types of sushi by $x_{1},\dots,x_{10}$. For computational efficiency, we consider only the partial rankings of the humans in our sushi dataset over the items $x_{1},\ldots,x_{5}$, treating these partial rankings as their ground-truth preferences. Each ranking is also associated with the fraction of the population in the dataset who had this preference over the types of sushi. A detailed breakdown of the most frequent sushi rankings is presented in Table˜3 (Appendix˜D).
Note that in this dataset, it is inherently impossible to distinguish between a) population-level heterogeneity in preferences, and b) user-level noise in reporting those preferences. To our knowledge, there is not a preference dataset that cleanly displays both a) and b). As such, in this section we make the assumption that users have reported their sushi preferences perfectly (no errors from b) and that the preference dataset simply reflects user heterogeneity (only difference is from a). In order to model user-level noise, we add noise (e.g., Mallows) on top of the ground-truth preferences within (kamishima2003nantonac). Furthermore, because this dataset does not include explicit human valuation scores for the sushi items, we assume item values within each ranking is determined by their Borda counts. Formally, given a ranking $\pi$ over $m$ items $\{x_{1},\ldots,x_{m}\}$, where $\pi(i)$ denotes the position of item $x_{i}$ (with $\pi(i)=1$ being the most preferred), the Borda score assigned to item $x_{i}$ is
@@ 8.1 Model Extensions
Other extensions could change our model of human-algorithm interaction. For example, our model of subset curation assumes that the human observes the set of presented items as an unordered set and is not biased by any inherent ordering of items. This is almost certainly not true in practice, as it has been demonstrated that humans are often biased by the order in which items are presented (e.g., see (joachims2007evaluating; agichtein2006improving; craswell2008experimental)), and this effect has been studied in human-algorithm collaboration more specifically (e.g. (mclaughlin2024designing; donahue2024listsbetteronebenefits; wang2022modeling)). It could also be possible that the human and algorithm undergo a sequence of back-and-forth interactions which ultimately end up at a final solution (as in (collina2025collaborative)), which may lead to different implications on personalization. Finally, our results on optimizing the algorithm for different social welfare functions assume that the algorithm has perfect knowledge of the distribution of human preferences. In practice, this is almost certainly not the case, and it could be useful to study settings where the algorithm must learn human preferences over multiple interactions.
@@ 8.2 Implications for Policy
In some real-life settings, humans have near-perfect ability to identify the “correct” outcome (e.g., the political opinion that they most closely identify with). However, in many real-life settings, humans may only be able to imperfectly (noisily) recover their own true preferences: prior research has shown that humans noisily pick the favorite item from sets, such as presented movies or items on a menu (agranov2017stochastic; plackett1975analysis; luce1959individual; hey1994investigating). In this setting, our results show that having algorithms that are “misaligned” in specific ways can be beneficial. This implies that designers of algorithms should maybe not always focus on designing algorithms that personalize to individual humans (eliminating misalignment): in settings where humans are noisy, personalizing may end up being harmful. This implies that algorithmic designers may want to take different strategies towards alignment depending on the context in which humans will use the tool, focusing on alignment more in noiseless settings and less in settings where humans are expected to know their own preferences less well.
@@
Our results also show that there can be benefits to algorithmic noise (e.g., positive temperature) when a single algorithm is serving a population of heterogeneous users. Benefits of randomization for goals related to fairness have been shown in other settings (e.g., (jain2024scarce; singh2018fairness)), though to our knowledge, this question is less well studied in the context of complementarity or similar benchmarks. Most generative AI tools have a constant temperature across multiple queries: our work suggests that for certain welfare objectives, it could be helpful to have temperature change dynamically, being lower in settings where human opinions are roughly in agreement and larger in settings where human opinions are more diverse.
@@ Acknowledgements
The research of Bhaskar Ray Chaudhury and Jiaxin Song was supported by NSF CAREER Grant CCF-2441580. The research of Kate Donahue was supported by the MIT METEOR Fellowship. The research of Parnian Shahkar was supported by NSF Grant CCF-2454115. We thank Kira Goldner (Boston University), Michael Kearns (University of Pennsylvania), Ariel Procaccia (Harvard University), Aaron Roth (University of Pennsylvania), Jann Spiess (Stanford), Kostas Kollias (Google Research), Manish Raghavan (MIT), Biaoshuai Tao (SJTU), Ming Yin (Purdue), as well as attendees at the Themis Science Talk at Amazon, the Human-AI workshop at EC’25, the Human-AI Complementarity for Decision Making Workshop at CMU, the AI and Science Group at Meta, and the Boston Economics and Computing Hub Workshops, for their valuable discussions and insightful suggestions. We also thank the anonymous reviewers for their constructive feedback.
@@ A.1 Basic Properties
Let $\mathfrak{S}(M)$ be the set of all rankings of $M$. Denote by $x_{i}\succ_{\pi}S$ for a set of items $S$ if $x_{i}$ is before any item $x_{j}\in S\setminus\{x_{i}\}$. A swap $(x_{i}x_{j})$ is valid with respect to $\pi$ and a ground-truth ranking $\pi^{*}$ if $\pi$ and $\pi^{*}$ differ on the relative ranking of $x_{i}$ and $x_{j}$. By the definition of inversion-monotonicity, if $(x_{i}x_{j})$ is a valid swap with respect to any given ranking $\pi$ and a ground-truth ranking $\pi^{*}$, then $\pi\circ(x_{i},x_{j})$ has a higher probability than $\pi$. Moreover, according to [donahue2024listsbetteronebenefits, Lemma 1], the number of inversions reduces by at least one after the swap.
@@ A.2 Properties of Mallows Model
$x_{i}$ ranked the first in $\pi$ $\frac{\exp(-\phi\cdot(i-1))}{Z_{m}(\phi)}$ [awasthi2014learning]
$x_{1}$ ranked at the $i$-th location in $\pi$ $\frac{\exp(-\phi\cdot(i-1))}{Z_{m}(\phi)}$ [awasthi2014learning]
$x_{1}$ ranked at the first $k$ locations in $\pi$ $\frac{Z_{k}(\phi)}{Z_{m}(\phi)}$ [awasthi2014learning]
$x_{i}$ ranked before $x_{j}$ with $i<j$ and k=ji+1k=j-i+1 in $\pi$ $\frac{k}{1-\exp(-\phi\cdot k)}-\frac{k-1}{1-\exp(-\phi\cdot(k-1))}$ [boehmer2023properties]
@@ A.2.1 Item Location
Then applying Equation˜3, since the sum of $\exp(-\phi\cdot\mathrm{inv}(\pi_{S}))$ over all permutations of $S$ is equal to $\prod_{i=1}^{k}Z_{i}(\phi)$ [mallows1957non], we then have
@@ A.2.2 Itemwise Comparison
Lemma 12 (Proposition 3.8 of [boehmer2023properties]).
@@
We first prove it holds for the case of $i=i^{\prime}$. Hence, by assumption, $j^{\prime}\geq j$. For any ranking $\pi$ that ranks $x_{i}$ before $x_{j}$ while ranking $x_{i}$ after $x_{j^{\prime}}$, we map it to a new one by swapping $x_{j}$ and $x_{j^{\prime}}$. By [donahue2024listsbetteronebenefits, Lemma 1], the number of inversions reduces by at least one, and the probability of the new ranking is weakly larger than the original one. Therefore, it implies that the probability of $x_{i}$ being before $x_{j}$ is smaller. By Lemma˜12, the probability of pairwise comparison only depends on the difference of the two indices. Therefore, the lemma holds for every $i,j,i^{\prime}$, and $j^{\prime}$. ∎
@@
We partition the set of permutations into $k$ sets based on the relative position of $x_{i(1)}$. Let $\mathfrak{S}_{t}$ be the set of permutations where $x_{i(1)}$ is placed at the $t$-th relative position among all items of $\mathcal{S}$. For any $\pi\in\mathfrak{S}_{t+1}$ with $t\leq k-1$, we map it to a permutation $\pi\in\mathfrak{S}_{t}$ by swapping $x_{i(1)}$ and the item that is placed at the $t$-th position among all items of $\mathcal{S}$. It is a valid swapping with respect to $\pi$ and $\pi^{*}$ by definition. By [donahue2024listsbetteronebenefits, Lemma 1], the probability of $\pi^{\prime}$ is at least $\exp(\phi)$ times the probability of $\pi$. By summing the inequality up over all permutations of $\mathfrak{S}_{t+1}$, we can conclude that $\operatorname*{\mathbb{P}}[\pi\in\mathfrak{S}_{t}]\geq\exp(\phi)\cdot\operatorname*{\mathbb{P}}[\pi\in\mathfrak{S}_{t+1}]$. Hence, by multiplying these inequalities, we can have
@@ C.2 Mixed Integer Linear Programming Formulation
Finally, we adapt the MIP of desir2016assortment and formulate the problem of finding the welfare-maximizing strategy as an MIP. desir2016assortment studies the assortment optimization problem under the Mallows model, where each item is associated with a nonnegative value $r_{i}$ and the goal is to pick a subset of items $\mathcal{S}$ to maximize the weighted sum of choice probabilities. Formally, the problem can be expressed as follows:
@@
desir2016assortment shows the choice probability of $x_{i}$ being the first among all items in $\mathcal{S}$ can be calculated by a dynamic programming algorithm. However, it is worth noting that the algorithm relies on prior knowledge that some item is known to be included in $\mathcal{S}$, which is further used as a “guess” of the MIP formulation. We next provide a dynamic programming of the choice probability with marginal change to the original algorithm of desir2016assortment without any prior knowledge.
Let $W(i,s,t)$ be the probability of item $x_{i}$ being chosen at position $s$ after $t$ steps of repeatedly insertion, where $i\in[m]$, $s\in[t]$, and $t\in[m]$. The following algorithm assumes the ground-truth ranking $\pi^{*}=(x_{1},\ldots,x_{m})$ without loss of generality. We use the same notations for constants $p_{t+1,s}$, $\gamma_{t+1,s}$ as in desir2016assortment. Probability $p_{t+1,s}$ is the probability of inserting item $x_{t+1}$ at position $s$ into a partial permutation of $x_{1},\ldots,x_{t}$. Also, $\gamma_{t,s}=\sum_{\ell=1}^{s}p_{t,\ell}$. Both sets of constants have closed-form expressions once the ground-truth ranking and the accuracy parameter are given.
@@
Compared to the original dynamic programming of desir2016assortment, we do not require a precondition that an item $x_{i}$ needs to be included in $\mathcal{S}$. Hence, we update the state space starting from step $t=1$ rather than $t=2$. As a result, if $x_{t}\in\mathcal{S}$, at the $t$-th step, there are two possible cases for item $x_{t}$ being selected at position $s$: 1) none of items $x_{1},\ldots,x_{t-1}$ is included in $\mathcal{S}$; therefore, item $x_{t}$ is selected at position $s$ with probability $p_{t,s}$; 2) at least one of items $x_{1},\ldots,x_{t-1}$ is included in $\mathcal{S}$; therefore, item $x_{t}$ is chosen only if all items of $x_{1},\ldots,x_{t-1}$ that are included in $\mathcal{S}$ are ranked after $x_{t}$ and $x_{t}$ is inserted at position $s$ (the same argument of desir2016assortment). Otherwise, if $x_{t}\notin\mathcal{S}$, the same argument of desir2016assortment still applies for updating the probabilities of choosing other items. ∎
@@ C.3 Uplift under Special Cases
Fix a set $S$ of size $k$. Notice that, when $x_{j}\notin S$, $\operatorname*{\mathbb{P}}[x_{H}=x_{j}\mid\rho[:k]=\mathcal{S}]=0$. Next, we consider the sum of the conditional probabilities for $x_{j}\in S$. We first show that, conditioned on $\rho[:k]=\mathcal{S}$, the partial ordering of the first $k$ items (denoted by $\operatorname{\rho}_{|\mathcal{S}}$) forms a Mallows distribution of items of $\mathcal{S}$. Observe that the number of inversions between $S$ and $M\setminus S$ for every permutation $\operatorname{\rho}$ such that $\rho[:k]=\mathcal{S}$ since the two sets of items are isolated. Hence, among all the permutations such that $\rho[:k]=\mathcal{S}$, the probability of each of them is proportional to $\exp(-\phi)$ raised to the power of the number of inversions inside $\mathcal{S}$, which implies that the partial ranking $\pi_{|S}$ forms a Mallows distribution of items of $\mathcal{S}$. Denote the $s$ items by $\mathcal{S}=\{x_{i(1)},\ldots,x_{i(k)}\}$ with $i(1)<\ldots<i(k)$. By the known property of Mallows model [awasthi2014learning, Lemma 2.5], the probability of item $x_{i(j)}$ being the first is equal to $\exp(-\phi\cdot(j-1))/Z_{k}(\phi)$. Summing it up for $j$ from $1$ to $i$, we get the probability of $x_{H}$ being one of $x_{1}$ to $x_{i}$ is equal to $Z_{s}(\phi)/Z_{k}(\phi)$, where $s$ is the size of the intersection of $\{x_{1},\ldots,x_{i}\}$ and $\mathcal{S}$. Therefore, the probability of the human picking one of the first $i$ items is equal to